H2 + I2 = 2HI (an elementary reaction)...If the volume of the container containing the gaseous mixture is increased to 2 times ,then final rate of reaction 
!- become 4 times
2 become ¼ th of original rate
3-become 2 times 
4 become ½ of the original rate.

Grade:12

3 Answers

Ramreddy IIT Patna
9 years ago
rate law = k[H2][I2]
when we increase the volume by 2 times the concentration becomes half. The rate of reaction becomes one fourth of the initial value. so, option (2) is correct.
….….….….….….….….….….….….….….….….….….….….….….….….….….…..
Naveen Kumar
9 years ago
H2 + I2 = 2HI
As it is elementary reaction, so order=molecularity=2
rate(r1)=k[H2][I2]
Now the volume is increased two times, so their concentration would be ½ of original value.
so rate(r2)=k*1/2*1/2*[H2]*[I2]=1/4*r1
.......................................................
Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

H2 + I2 --> 2HI elementary reaction,
order = molecularity = 2
rate (r1) = k[H2][I2]
Now the volume is increased two times, so their concentration would be ½ of original value.
so rate(r2) = k*1/2*[H2]*1/2*[I2] = 1/4*r1

Thanks and Regards